OpenStudy (anonymous):

Type the equation of the given line in standard form. The line with m = 4 and the point (0, -9).

OpenStudy (ajprincess):

The equation of a line passing through a poijt \((x_1,y_1)\) and with the slope m is \(\large\frac{y-y_1}{x-x_1}=m\) Can u find nw?

OpenStudy (anonymous):

So that would be... (y+9)/(x-0)=4 (y+9)/x=4 But what do I do from there? Do I cross multiply?

OpenStudy (anonymous):

Nevermind, I got it :P Thank you!

OpenStudy (ajprincess):

yw.